OpenStudy (anonymous):

Write an equation of the line whose slope is 3 and which passes through the point (1,6).

OpenStudy (anonymous):

Is it y=mx+b form?

OpenStudy (lieutenantgeneral):

>?

OpenStudy (anonymous):

I got y=3x+3

OpenStudy (anonymous):

It's an open ended question

OpenStudy (lieutenantgeneral):

2y = 6x + 12 gives y = 3x + 6, so m = 3

OpenStudy (anonymous):

so, my answer is wrong?

OpenStudy (anonymous):

@LieutenantGeneral

OpenStudy (lieutenantgeneral):

no

OpenStudy (anonymous):

I didn't read everything you wrote

OpenStudy (lieutenantgeneral):

lol r u girl or boy?

OpenStudy (anonymous):

a girl lol

OpenStudy (anonymous):

SO, for 2y=6x+12, you divide by 2?

OpenStudy (the_fizicx99):

You have the slope and 2 points, y-y1=m(x-x1) y-6 = 3(x-1) y-6 = 3x - 3 +6 6 y = 3x + 3